OpenStudy (anonymous):

After graduation, B. Kind pledges to make an annual donation to his college. The first year he will donate $200. Every year after that he will increase his donation by 12%. How much will Mr. Kind be giving to the college over the next 20 years? This problem would require formulas for: D) geometric series B) arithmetic series C) geometric sequences A) arithmetic sequences

OpenStudy (anonymous):

I believe this would be a geometric sequence

OpenStudy (anonymous):

see it will go like this 1st year = $200 2nd year =$200+(0.12*$200)=$200+{2-1}(0.12*$200) 3rd year=$200+(0.12*$200)+(0.12*$200)=$200+2*(0.12*$200)=$200+{3-1}(0.12*$200) 4th year=$200+{4-1}(0.12*$200).......so on to 20th year=$200+{20-1}(0.12*$200) hence it resemble a20=a+(20-1)d where a is 1st term,d is common difference and a20 is 20th term of the series this is the general formula of AP series.

OpenStudy (anonymous):

So its geometric sequence?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

@kelliegirl33 and @sweetpink365 its a arithmetic series

OpenStudy (anonymous):

Oh ok thank you :)

OpenStudy (anonymous):

you sure....because you are multiplying to find the difference

OpenStudy (anonymous):

@kelliegirl33 if it would be a geometric series then we would see term like this $200(0.12*200)^19

OpenStudy (anonymous):

this is a geometric series http://www.regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm

OpenStudy (anonymous):

@kelliegirl33 to be frank i m not sure about GP but what the question ask and what I did its an AP

OpenStudy (anonymous):

ok....I am probably wrong....thanks for explaining